OpenStudy (anonymous):

can anyone help ??? Which of the following is not a way to represent the solution of the inequality 7 - 9x - (x + 12) <=25? x >=-3 x <=-3 -3 <=x A number line with a closed circle on negative 3 and shading to the right.

OpenStudy (anonymous):

x <=-3

OpenStudy (anonymous):

please explain

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you have this 7 - 9x - (x + 12) <=25 7-9x-x-12<=25 -10x-5<=25 -10x <= 30 x >= -3 now this is the correct answer but you are looking for the incorrect choice thus the only incorrect choice is x<=-3

OpenStudy (anonymous):

can u help with another ?

OpenStudy (anonymous):

post it

OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality y < 2x + 6 Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin.

OpenStudy (anonymous):

can you guess the choice now??

OpenStudy (anonymous):

according to the site, the answer is line. -_- my origional guess was C

OpenStudy (anonymous):

@rizwan_uet